when 247 is divided by 6, the remainder is r, and when 247 is divided by 12,the remainder is s. what is the value of r – s ?
a -6 b -1 c 0 d 1 e 6

Answer by Edwin McCravy(20081)

when 247 is divided by 6,

   41
6)247
  24 
    7
    6
    1 

the remainder is r,

I got a remander of 1, so r is 1

and when 247 is divided by 12,

    20
12)247
   24
     7
     0
     7

the remainder is s.

I got a remainder of 7, so s is 7

what is the value of r – s ?

Since r is 1, and s is 7,  then r - s = 1 - 7 = -6

So the answer is -6, which is choice (a)

Edwin